# Monotonic Path-Specific Effects: Application to Estimating Educational Returns

### v1

@@ 5 Results without imputation of missing covariates
In the main text, I report estimates of the MPSE decomposition for the ATE of high school attendance on logged annual earnings for the full sample of NLSY97 respondents with non-missing educational information and non-missing earnings ($N=7,305$). A very large number (approximately $50\%$) of these respondents are missing information on one or more of the covariates $(X,Z)$ used in the models in order to identify the decomposition components, due especially to the large amount of missingness on variables such as parental assets ($25\%$ of the sample), and the Armed Services Vocational Aptitude Battery (ASVAB) test ($20\%$ of the sample). To assess the sensitivity of my primary conclusions to the use of multiple imputation as opposed to dropping observations with missing values, I replicate my DML and RWR estimates on a non-imputed analytic sample ($N=3,735)$. Figure 6 below shows the results of this exercise. For both estimation procedures, results under multiple imputation and non-imputation are highly similar. As is to be expected, imputation reduces standard errors significantly, especially for the parametric RWR procedure. Further, the greatest variability between imputed and non-imputed results come from effects pertaining to high school completion, perhaps because patterns of missingness are correlated with educational attainment. Despite this, because the total effect $\tau_{0}$ and direct effect $\theta_{0}$ are similarly attenuated in the imputed sample, the overall conclusion about the importance of the direct effect in explaining the ATE remains unaffected.
@@
6 Results under alternative definitions of earnings
In the main text, I report estimates of the MPSE decomposition components for the ATE of high school attendance on logged annual earnings. Logged annual earnings in the main text are defined as the log of observed annual earnings plus a small constant of \$1,000\\$1,000, in order to accommodate respondents with zero observed annual earnings. In order to assess the sensitivity of the reported results to the choice of this constant, I replicate the main analyses under alternative definitions of earnings. Figure 7 reports estimates of the direct and indirect effects under a series of different constants $c$ added to pre-logged annual earnings, for $c\in\{10,100,1000\}$, while Figure 8 shows estimates of these direct and indirect effects for observed annual earnings in dollar values. Beginning with Figure 7, I see that, while for the indirect effects $\theta_{1}$, $\theta_{2},$ and $\theta_{3}$, both DML and RWR estimates are quite consistent under these different constants, estimates of the total effect $\tau_{0}$ as well as the direct effect $\Delta_{0}$ are quite sensitive to the choice of constant. Specifically, lower constant values correspond with large increases in the DML estimate of $\tau_{0}$ from $0.67$ ($c=1000$) to $1.20$ ($c=10$), and of $\Delta_{0}$ from $0.47$ ($c=1000$) to $0.89$ ($c=10$). This is because individuals with less than a high school degree are more likely than their higher-educated counterparts to have zero or low earnings, making their logged earnings rather sensitive to the choice of constant. Nevertheless, because the total effect $\tau_{0}$ and direct effect $\theta_{0}$ are similarly affected by the change in constant value, the importance of the direct effect in explaining the ATE of high school completion is reinforced. In particular, the proportion of the total effect that is direct is estimated to be $70\%$, $74\%$ and $72\%$ under each of $c=10,100,1000$. Turning next to Figure 7, under DML, estimates of high school completion increases earnings in expectation by roughly \$16,500\\$16,500, corresponding to an earnings return of approximately $53$ relative to a baseline of \$31,300\\$31,300 without high school graduation ($\mathbb{E}[Y(0)]$) . Almost half ($\frac{7824}{16454.}\cdot 100=47.5\%$) of the total effect is estimated to operate directly, with $29\%$ and $22\%$ mediated via college attendance and college completion, respectively.
@@ 6 Results under alternative definitions of earnings
7 Extension to multivalued, discrete mediators
The main text considers a decomposition of the ATE in the case of binary monotonic mediators (i.e., educational transitions), but the framework can straightforwardly be extended to accommodate categorical transitions. Such a decomposition is especially appealing when we consider the variegated and complex trajectories that individuals take through the US postsecondary system, that dichotomizing transitions invariably misses. In particular, in the early 2010s, under $40\%$ of high school graduates immediately enrolled in a four-year college, with around $30\%$ immediately attending a two-year college. Approximately one third of immediate two-year college attendees will then progress to a four-year college; indeed, nearly $50\%$ of BA recipients previously attended a public two-year college in their educational careers. To illustrate the MPSE decomposition for multivalued, discrete mediators in the case of a single mediator, consider high school graduation $A$, and immediate college enrollment $M\in\{\text{none},\text{2yr},\text{4yr}\}$ denotes whether an individual did not pursue postsecondary education, or instead attended a 2-year or 4-year college. Assumption 2 is similarly assumed to hold in this context, such that $M(0)=0$ (high school non-completers cannot pursue any form of postsecondary education). Under this assumption, I obtain the following multivariate decomposition:
ATE $=\mathbb{E}[Y(1)-Y(0)]$
$=\mathbb{E}[Y(1,0)-Y(0)]$
$+\mathbb{P}[M(1)=\text{2yr}]\mathbb{E}[Y(1,\text{2yr})-Y(1,0)]+\text{cov}\big{[}\mathbb{I}[M_{1}(1)=\text{2yr}],Y(1,\text{2yr})-Y(1,0)\big{]}$ (14)
$+\mathbb{P}[M(1)=\text{4yr}]\mathbb{E}[Y(1,\text{4yr})-Y(1,0)]+\text{cov}\big{[}\mathbb{I}[M_{1}(1)=\text{4yr}],Y(1,\text{4yr})-Y(1,0)\big{]}.$ (15)
Each of these quantities can be identified under Assumptions 3-5 above. Further, Equations 14 and 15 can be further decomposed into the direct and gross effects of each of these transitions in a way analogous to Equation 4 (for example, the gross effect of 2-year attendance can be decomposed into its direct effects, net of subsequent transfer to a 4-year college, and its indirect effect via 4-year transfer, which itself can be further decomposed into direct and gross effects).
8 Description of EIFs used in empirical illustration
@@
9 Proofs and technical details
9.1 EIFs for $\eta_{k}$ and $\theta_{k}$ terms (Proposition 1)
@@
9.2 Semiparametric efficiency (Proposition 2)
@@
where the penultimate equality follows from Proposition 1, and the final equality follows from Lemma 1.
@@ 10 Derivation of RWR procedures
11 Derivation of bias formulae for sensitivity analysis

### v2

@@ E Results without imputation of missing covariates
To assess the sensitivity of my primary conclusions to these issues, I undertake two exercises. First, I replicate my DML and RWR estimates on a non-imputed analytic sample, dropping observations with missing values ($N=3,735)$. Figure 8 below shows the results of this exercise. For both estimation procedures, results under multiple imputation and non-imputation are highly similar. As is to be expected, imputation reduces standard errors significantly, especially for the parametric RWR procedure. Further, the greatest variability between imputed and non-imputed results come from effects pertaining to high school completion, perhaps because patterns of missingness are correlated with educational attainment. Despite this, because the total effect $\tau_{0}$ and direct effect $\theta_{0}$ are similarly attenuated in the imputed sample, the overall conclusion about the importance of the direct effect in explaining the ATE remains unaffected.
Second, to assess whether missingness in field of study reflects meaningful differences in educational progression versus measurement limitations, I re-estimated the entire MPSE decomposition excluding the STEM indicator from the intermediate confounder set $Z$. The resulting estimates (presented in Table 6) are nearly identical to those reported in the main text, indicating that the main empirical findings are not sensitive to whether field of study is included in the intermediate confounder set. Nevertheless, the question of whether major declaration itself constitutes an important intermediate transition (prior to BA completion) is an important direction for future research.
Table 4: Proportion of missing background covariates by educational level
Parental income
Parental education
Household net worth
ASVAB percentile
High school GPA
Peers’ expectations (75th)
Peers’ expectations (90th)
Property stolen at school
HS Non-Completers 1144 0.058 0.069 0.233 0.302 0.097 0.021 0.021 0.034
HS Graduates 3443 0.033 0.036 0.238 0.203 0.008 0.015 0.015 0.011
College Goers 885 0.040 0.030 0.270 0.159 0.016 0.006 0.006 0.005
BA Completers 1246 0.019 0.023 0.255 0.127 0.009 0.007 0.007 0.006
Grad. School Goers 587 0.029 0.020 0.307 0.148 0.019 0.007 0.007 0.005
Table 5: Proportion of missing intermediate covariates (transcript-based) by educational level
NN
STEM major
College GPA
HS Non-Completers 1144 1.000 1.000
HS Graduates 3443 1.000 1.000
College Goers 885 0.103 0.357
BA Completers 1246 0.015 0.272
Grad. School Goers 587 0.019 0.274
Table 6: Direct Effects ($\Delta_{k}$), Probabilities ($\pi_{k}$) and Covariance Terms ($\eta_{k}$) Involved in Decomposition via Debiased Machine-Learning (DML), without STEM degree.
DML 0.462 0.197 0.463 0.124 0.427 0.554 0.313 0.006 0.007 -0.016
(0.059) (0.034) (0.046) (0.029) (0.009) (0.014) (0.022) (0.007) (0.007) (0.009)
@@
F Results under alternative definitions of earnings
In the main text, I report estimates of the MPSE decomposition components for the ATE of high school attendance on logged annual earnings. Logged annual earnings in the main text are defined as the log of observed annual earnings plus a small constant of $\mathdollar 1,000$, in order to accommodate respondents with zero observed annual earnings. In order to assess the sensitivity of the reported results to the choice of this constant, I replicate the main analyses under alternative definitions of earnings. Figure 9 reports estimates of the direct and indirect effects under a series of different constants $c$ added to pre-logged annual earnings, for $c\in\{10,100,1000\}$, while Figure 10 shows estimates of these direct and indirect effects for observed annual earnings in dollar values. Beginning with Figure 9, We see that, while for the indirect effects $\theta_{1}$, $\theta_{2},$ and $\theta_{3}$, both DML and RWR estimates are quite consistent under these different constants, estimates of the total effect $\tau_{0}$ as well as the direct effect $\Delta_{0}$ are quite sensitive to the choice of constant. Specifically, lower constant values correspond with large increases in the DML estimate of $\tau_{0}$ from $0.67$ ($c=1000$) to $1.20$ ($c=10$), and of $\Delta_{0}$ from $0.47$ ($c=1000$) to $0.89$ ($c=10$). This is because individuals with less than a high school degree are more likely than their higher-educated counterparts to have zero or low earnings, making their logged earnings rather sensitive to the choice of constant. Nevertheless, because the total effect $\tau_{0}$ and direct effect $\theta_{0}$ are similarly affected by the change in constant value, the importance of the direct effect in explaining the ATE of high school completion is reinforced. In particular, the proportion of the total effect that is direct is estimated to be $70\%$, $74\%$ and $72\%$ under each of $c=10,100,1000$. Turning next to Figure 9, under DML, estimates of high school completion increases earnings in expectation by roughly $\mathdollar 16,500$, corresponding to an earnings return of approximately $53$ relative to a baseline of $\mathdollar 31,300$ without high school graduation ($\mathbb{E}[Y(0)]$). Almost half ($\frac{7824}{16454.}\cdot 100=47.5\%$) of the total effect is estimated to operate directly, with $29\%$ and $22\%$ mediated via college attendance and college completion, respectively.
@@ F Results under alternative definitions of earnings
G Results for the NLSY79 cohort
Because NLSY97 respondents are observed only through age 36, the main analysis uses logged average earnings over ages 32–36. To evaluate whether later-life earnings alter the decomposition—especially for graduate schooling, whose returns may rise after early adulthood—I also report results using the NLSY79 cohort, for whom we can observe earnings up until age 44.
All variables are defined analogously, although several differences between the two surveys are of note. First, the NLSY79 contains a narrower set of baseline covariates. Whereas the NLSY97 includes high school GPA and a rich set of school characteristics (disciplinary environment, parental assets, school safety, and peer-context measures), the NLSY79 does not. Consequently, the covariate vector in this replication consists of a more limited set of demographic and family background characteristics: indicators of gender and race/ethnicity; parental socioeconomic status (income, educational attainment, and occupation); family structure and sibship size; a cognitive ability measure based on the Armed Forces Qualification Test (AFQT); and the respondent’s expectations for their own educational attainment. The NLSY79 also provides several proxies for cultural resources in the home at age 14—specifically, whether the household regularly received magazines or newspapers and whether any household member held a library card. These measures capture elements of the early educational environment but are considerably narrower than the school-context and academic performance variables available in the NLSY97.
Second, the NLSY79 lacks transcript-based measures of college major or GPA, so there are no intermediate covariates $Z$ between the BA and graduate-school transitions. Identification of the MPSE components therefore relies on stronger assumptions than in the NLSY97 analysis, as academic performance and field-of-study sorting cannot be adjusted for when isolating the contribution of BA and post-BA educational pathways.
Figures 11 and 12 report estimates of the direct and indirect components of the MPSE decomposition for the NLSY79 cohort, for earnings at age 32-26 (top panel) and at age 35-44 (bottom panel). Figure 11 presents results when a constant $c\in\{10,100,1000\}$ is added to annual earnings prior to logging, while Figure 12 shows the decomposition for observed earnings in dollar values. Table 7 also summarizes the estimated direct effects ($\Delta_{k}$), stage-specific transition probabilities ($\pi_{k}$), and covariance components ($\eta_{k}$) for the NLSY79 cohort for earnings at age 32-26 and at age 35-44.  [^11] There are several differences in the estimated effects in the NLSY79 versus the NLSY97 cohorts. Notably, across specifications, the total effect of high school completion is smaller in the NLSY79 cohort ($\tau_{0}=0.4$ for log(Y+1000)(Y+1000), compared with $\tau_{0}=0.675$ in the NLSY97), consistent with lower marginal returns to schooling for the earlier cohort. The contribution of college attendance to the total effect is also estimated to be negligible ($\theta_{1}=0.03$) compared with in the NLSY79 cohort ($\theta_{1}=0.092$), which may reflect the fact that mediation via “college attendance” in this decomposition captures any postsecondary enrollment, as opposed to mediation via 4-year attendance in the NLSY97, as the NLSY79 public use dataset does not distinguish clearly between two-year and four-year colleges.
Within the NLSY79 cohort, across the earlier and later earnings windows, results are remarkably stable for the log-transformed specifications. As shown in Figure 11, estimates of $\tau_{0}$ and $\theta_{k}$ are nearly indistinguishable across log(Y+10)(Y+10), log(Y+100)(Y+100), and log(Y+1000)(Y+1000), with only mild divergence when the largest constant is applied. By contrast, decompositions based on raw earnings (Figure 12) exhibit larger differences across age windows. The larger total effect $\tau_{0}$ at ages 35–44 than at ages 32–36, for example, results from the fact that the same proportional earnings premium associated with high school completion translates into a larger dollar difference as overall earnings rise with age. Even so, the ordering and relative magnitudes of the direct and indirect pathways remain consistent across specifications.
Importantly, the graduate-school pathway contributes almost identically to the total effect at ages 32–36 and 35–44, even though one might reasonably expect its contribution to increase at later ages as the graduate-school earnings premium becomes larger.
Table 7 clarifies the source of this stability. The continuation effect ($\Delta_{3}$) nearly doubles between the two windows, but the covariance component $\eta_{3}$ declines, indicating a weaker alignment between the propensity to attend graduate school and the incremental graduate-school earnings premium at later ages. Early in adulthood, individuals most likely to pursue graduate education tend to experience the largest immediate gains from doing so, generating a larger covariance term. By ages 35–44, graduate-school earnings advantages are more broadly distributed across degree holders rather than being concentrated among those with the highest propensity to attend. Because the overall probability of traversing the entire trajectory from high school completion to graduate school remains small, these offsetting forces produce a nearly unchanged $\theta_{3}$ across age windows.
Figure 11: Decomposition of the Average Total Effect (ATE) of high school graduation on logged earnings under alternative definitions of earnings. The figure shows estimates of the total effect ($\tau_{0}$) as well as the continuation effects $\Delta_{0},\Delta_{1},\dots,\Delta_{K}$ when constants of 10, 100, and 1000 are added to annual earnings (in dollars) prior to taking the log.
Figure 12: Decomposition of the Average Total Effect (ATE) of high school graduation on observed annual earnings in dollars (no transformation). The figure shows estimates of the total effect ($\tau_{0}$) as well as the continuation effects $\Delta_{0},\Delta_{1},\dots,\Delta_{K}$ using raw annual earnings.
Table 7: Direct Effects ($\Delta_{k}$), Probabilities ($\pi_{k}$) and Covariance Terms ($\eta_{k}$) Involved in Decomposition via Debiased Machine-Learning (DML) for NLSY79 cohort.
Earnings Age 32–36 0.300 0.070 0.520 0.150 0.480 0.330 0.120 -0.010 -0.040 0.020
Earnings Age 35–44 0.330 0.140 0.440 0.290 0.480 0.330 0.120 -0.010 -0.030 -0.010
(0.070) (0.040) (0.040) (0.050) (0.010) (0.010) (0.010) (0.010) (0.010) (0.010)
(0.060) (0.040) (0.040) (0.040) (0.010) (0.010) (0.010) (0.010) (0.010) (0.010)
H Extension to categorical mediators
The main text considers a decomposition of the ATE in the case of binary monotonic mediators (i.e., educational transitions), but the framework naturally extends to settings with categorical transitions. Such an extension is especially appealing in the context of U.S. higher education, where individuals follow increasingly variegated pathways. In the early 2010s, fewer than $40\%$ of high school graduates enrolled directly in a four-year college, whereas roughly $30\%$ entered a two-year college, and close to one third of these later transferred to a four-year program. Nearly half of all BA recipients nationally had attended a two-year college at some point in their educational careers. A similar situation arises at the transition to postgraduate education: “graduate school” encompasses multiple substantively distinct routes (e.g. Master’s degrees, professional programs, PhDs), each featuring different patterns of selection and labor-market returns. Dichotomizing such transitions discards precisely the heterogeneity that is often of substantive interest.
To illustrate the extension, consider a single intermediate transition $M_{k}$ that now takes values in a finite set of mutually exclusive and exhaustive categories,
$\mathcal{H}_{k}=\{h_{1},\ldots,h_{H}\},$
rather than a binary indicator. For each category $h\in\mathcal{H}_{k}$, let $M_{k}(h)$ denote the potential mediator value under category $h$, and let $Y(1_{k},h)$ denote the potential outcome under completion of all prior transitions $1_{k}=(A=1,M_{1}=1,\dots,M_{k-1}=1)$ with the mediator $M_{k}$ set to $h$.
The net effect of transition $k$ can then be written as
$\tau_{k}=\Delta_{k}\;+\;\sum_{h\in\mathcal{H}_{k+1}}\bigl(\pi_{{k+1},h}\,\Delta_{{k+1},h}\;+\;\eta_{{k+1},h}\bigr),$ (18)
where
$\Delta_{k,h}$ $\equiv\mathbb{E}[Y(1_{k},h)-Y(1_{k},h_{0})],\qquad h_{0}\in\mathcal{H}_{k}\ \text{(baseline)},$
$\pi_{k,h}$ $\equiv\mathbb{E}\!\left[\mathbf{1}\{M_{k}(1_{k})=h\}\right],$
$\eta_{k,h}$ $\equiv\operatorname{cov}[\mathbf{1}\{M_{k}(1_{k})=h\},\ Y(1_{k},h)-Y(1_{k},h_{0})].$
These terms generalize the binary case: $\Delta_{k,h}$ is the direct effect of category $h$ relative to $h_{0}$; $\pi_{k,h}$ is the counterfactual probability of attaining $h$ given completion of all prior transitions; and $\eta_{k,h}$ captures the alignment between the payoff $Y(1_{k},h)-Y(1_{k},h_{0})$ and the propensity to realize category $h$. The direct effects $\Delta_{k,h}$ and path probabilities $\pi_{k,h}$ remain identified under Assumptions 35. However, when $|\mathcal{H}_{k}|>2$, the covariance terms are no longer separately nonparametrically identified. In the binary case,
$\tau_{k}=\Delta_{k}+\pi_{k}\Delta_{k}+\eta_{k}$
contains a single $\eta_{k}$ and one identifying restriction. By contrast, (18) contains multiple unknowns $\{\eta_{k,h}\}$ but only one restriction, implying that the vector of covariance components cannot be uniquely recovered without additional structure—for example, assuming homogeneity across categories, proportionality, or a parametric model for effect heterogeneity.
Graduate school example (NLSY97)
In my empirical application, the $k=3$ transition corresponds to the education decision following BA completion. I disaggregate this transition into three categories:
$\mathcal{H}_{3}=\{\mathrm{BA\!-\!only},\;\mathrm{MA},\;\mathrm{PhD}\},$
with BA-only as the baseline $h_{0}$. Among BA completers in the NLSY97, roughly $73\%$ do not pursue any graduate degree, $22\%$ complete a Master’s degree, and only $5\%$ complete a professional or PhD degree.  [^12] Professional (DDS, MD and JD) and PhD degree holders are pooled due to small cell sizes. Stage-specific causal contrasts are:
$\Delta_{3,h}=\mathbb{E}\!\bigl[Y(1_{3},h)-Y(1_{3},\mathrm{BA\!-\!only})\bigr],\qquad h\in\{\mathrm{MA},\mathrm{PhD}\}.$
The continuation value that enters the full MPSE decomposition is
$\theta_{3}=\left(\prod_{j=1}^{2}\pi_{j}\right)\sum_{h\in\{\mathrm{MA},\mathrm{PhD}\}}\bigl(\pi_{3,h}\,\Delta_{3,h}+\eta_{3,h}\bigr).$ (19)
Figure 13 displays the resulting decomposition for MA and professional/PhD pathways, and Table 8 reports the estimated components. The continuation effects for the graduate-school transition are small for both pathways, with $\widehat{\theta}_{3,\mathrm{MA}}=0.0119$ and $\widehat{\theta}_{3,\mathrm{PhD}}=0.0077$, corresponding to earnings effects of high school graduation via these transitions of $1.2\%$ and $0.8\%$, respectively. Although the continuation effect is slightly lower for the PhD/professional pathway, this masks a much larger underlying causal contrast: the net effect of completing a PhD or professional postgraduate program relative to not pursuing postgraduate study is $\widehat{\Delta}_{3,\mathrm{PhD}}=0.521$ (a $68\%$ earnings premium), compared to a more modest $\widehat{\Delta}_{3,\mathrm{MA}}=0.213$ (a $24\%$ premium) for Master’s degrees. The similarity in the overall continuation effects instead reflects stark differences in the counterfactual probabilities of each pathway: only about $2\%$ of BA completers pursue a PhD-level degree ($\widehat{\pi}_{3,\mathrm{PhD}}\approx 0.021$), whereas Master’s attainment is roughly six times more common ($\widehat{\pi}_{3,\mathrm{MA}}\approx 0.134$).
Table 8: Estimated components for MA and PhD pathways in the graduate-school transition.
$\Delta_{3,MA}$ $\Delta_{3,PhD}$ $\pi_{3,MA}$ $\pi_{3,PhD}$ $\eta_{3,MA}$ $\eta_{3,PhD}$
DML 0.213 0.521 0.134 0.021 0.022 0.022
(0.032) (0.032) (0.007) (0.002) (0.008) (0.008)
Figure 13: Continuation effects for Master’s and Professioal/PhD pathways in the NLSY97.
I Description of EIFs used in empirical illustration
@@
J Proofs and technical details
J.1 EIFs for $\eta_{k}$ and $\theta_{k}$ terms (Proposition 3.1)
@@
J.2 Semiparametric efficiency (Proposition 3.2)
@@
where the penultimate equality follows from Proposition 3.1, and the final equality follows from Lemma J.1.
@@
L Derivation of bias formulae for sensitivity analysis